OpenStudy (anonymous):

Factor completely. 3x3 – 48x A. 3x(x – 4)(x – 4) B. 3(x + 4)(x – 4) C. 3x(x – 4)(x + 4) D. 3x(x – 16) @jagr2713

OpenStudy (anonymous):

@ken

OpenStudy (anonymous):

@KendrickLamar2014 @jagr2713

jagr2713 (jagr2713):

hat is the largest number that divides evenly into 3x3 and −48x?

jagr2713 (jagr2713):

what*

OpenStudy (anonymous):

3?

OpenStudy (rainbow_rocks03):

plz help

jagr2713 (jagr2713):

Good so What is the highest degree of x that divides evenly into 3x3 and −48x?

OpenStudy (anonymous):

help with what?? @rainbow_rocks03

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

\[3x^3-48x=3x\left( x^2-16 \right)=?\] \[a^2-b^2=\left( a+b \right)\left( a-b \right)\]

jagr2713 (jagr2713):

nope degree not number @Nin

OpenStudy (rainbow_rocks03):

help with question

OpenStudy (anonymous):

48?

OpenStudy (rainbow_rocks03):

need to understand how to get my answer

jagr2713 (jagr2713):

@rainbow_rocks03 please post your question in the ask a question bo

OpenStudy (rainbow_rocks03):

I did

jagr2713 (jagr2713):

nope its X b.c thats the degree so now we multiplying the results above, the GCF is 3x

OpenStudy (anonymous):

yes please, we are working on something!! @rainbow_rocks03

OpenStudy (anonymous):

okay!@jagr2713

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

So now we factor out the GCF so put the GCF as the first term and then in teh parentheses divide each term by the GCF \[3x (\frac{ 3x2 }{ 3x }+\frac{ -48x }{ 3x }\]

jagr2713 (jagr2713):

So now we Simplify each term in parentheses and get?

OpenStudy (anonymous):

so 2+ 16x?

jagr2713 (jagr2713):

nope we simplify each term in parentheses and get 3x(x2-16)

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

So this ik you remember so we have x2−16 fits the form a2−b2 Rewrite it where a=x and b=4 3x(x2−42)

jagr2713 (jagr2713):

So now use the difference of squares \[a ^{2}-b ^{2}=(a+b)(a-b)\]

jagr2713 (jagr2713):

Plug it in :D

OpenStudy (anonymous):

(x+4)(x-4)

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

dont forget the 3x so the answer is?

jagr2713 (jagr2713):

So we get? igtg soon

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@jazzyspazzy @Jaynator495 @kdenice123 @inowalst @KaylaIsBae

jagr2713 (jagr2713):

Nope i said 3x

OpenStudy (anonymous):

@jagr2713 c?

jagr2713 (jagr2713):

Correct :D

OpenStudy (anonymous):

i didnt see that it was just 3

jagr2713 (jagr2713):

Well time to go :D ~Have a great day and great job :D

OpenStudy (anonymous):

thank you b!